It was believed ( from past data) that professionals across a particular city had an average net income was Rs 54000 per month with a standard deviation of Rs12000. A researcher wanted to determine whether the net income has changed over a period of time. He took a survey of 100 professionals and found that the average net income is Rs 56000. Test at 1% level of significance.


1
Expert's answer
2021-02-24T07:28:11-0500

We have that

"\mu=54000"

"\sigma=12000"

"n=100"

"\bar x=56000"

"\alpha=0.01"

"H_0:\mu=54000"

"H_a:\mu \ne54000"

The hypothesis test is two-tailed.

The population standard deviation is known and the sample size is large (n≥30) so we use z-test.

The critical value for α=1% is Z0.01 = ±2.57

The critical region is Z < -2.57 and Z > 2.57

Test statistic:


"Z_{test}=\frac{\bar x -\mu}{\frac{\sigma}{\sqrt n}}=\frac{56000 -54000}{\frac{12000}{\sqrt {100}}}=1.66"

Since -2.57 < 1.64 < 2.57 thus the Ztest does not fall in the rejection region we fail to reject the null hypothesis. Therefore there is no sufficient evidence to conclude that the net income has changed over a period of time.